007 (part 1 of 2) 10.0 points A cylindrical fishing reel has a mass of 0.791 kg and a radius of 2.36 cm. A fric tion clutch in the reel exerts a restraining torque of 1.21 N - m if a fish pulls on the line The fisherman gets a bite, and the reel be- gins to spin with an angular acceleration of 65.1 rad/s2. How much line unwinds in 0.632 s? Answer in units of cm

How much line unwinds in 0.632 s?

I = 1/2 MR^2
Answer #1

angular acceleration , a = 65.1 rad/s^2

time , t = 0.632 s

Now, for the angle rotated by the reel

theta = 0.50 * at^2

theta = 0.50 * 65.1 * 0.632^2

theta = 13 radian

line unwinds = 13 * 0.0236

line unwinds = 0.31 m

the unwinding of line is 0.31 m